# A data-driven framework for team selection in Fantasy Premier League

### v2

@@ Abstract
Fantasy football is a billion-dollar industry with millions of participants. Under a fixed budget, managers select squads to maximize future Fantasy Premier League (FPL) points. This study formulates lineup selection as data-driven optimization and develops deterministic and robust mixed-integer linear programs that choose the starting eleven, bench, and captain under budget, formation, and club-quota constraints (maximum three players per club). The objective is parameterized by a hybrid scoring metric that combines realized FPL points with predictions from a linear regression model trained on match-performance features identified using exploratory data analysis techniques. The study benchmarks alternative objectives and cost estimators, including simple and recency-weighted averages, exponential smoothing, autoregressive integrated moving average (ARIMA), and Monte Carlo simulation. Experiments on the 2023/24 Premier League season show that ARIMA with a constrained budget and a rolling window yields the most consistent out-of-sample performance; weighted averages and Monte Carlo are also competitive. Robust variants improve some objectives but are not uniformly superior. The framework provides transparent decision support for fantasy roster construction and extends to FPL chips, multi-week rolling-horizon transfer planning, and week-by-week dynamic captaincy.
keywords:
data science and management , fantasy football , data-driven optimization , time series forecasting , machine learning
[…]
@@ 1 Introduction
Against this backdrop, fantasy football has attracted attention across research and practitioner communities. Qualitative studies examine impacts on related markets and behavior, including merchandise sales [drayer2010effects], game attendance [nesbit2010impact], and participant well-being [wilkins2021exploring]. Quantitative research has explored performance prediction and strategy design using econometric and machine learning techniques. The increasing availability of public data and analytics tooling has further lowered barriers to empirical work and large-scale evaluation in this domain.
Despite this progress, most quantitative FPL studies emphasize forecasting weekly player points, with comparatively less attention to integrating those forecasts into a decision model that captures the full set of FPL constraints and jointly selects the starting XI and captain. Decision-analytic approaches exist [matthews2012competing], but captaincy is typically handled outside the optimization pipeline (e.g., via crowds or heuristics) rather than optimized jointly [bhatt2019should]. Moreover, robustness to estimation error is rarely addressed, and there is limited comparative evidence on how alternative objective functions, selection budget, or cost-vector estimation methods affect out-of-sample outcomes within a unified optimization framework.
@@
  • We formulate FPL lineup selection as a data-driven optimization problem and develop deterministic and robust integer models that choose the optimal starting XI, bench players, and the captain under budget, formation, and club-quota constraints.
  • We propose a hybrid scoring metric that combines realized FPL points with predictions from a ridge regression model trained on explainable match-performance features.
  • We build a unified pipeline that benchmarks alternative objectives and multiple forecasting approaches (including averaging techniques, exponential-smoothing, ARIMA, and simulation-based estimators) and provide an empirical assessment on the 2023/24 Premier League season with managerial insights into emergent formations and player selection patterns. We also implemented a rolling window approach and analyzed how constraining the budget for the starting XI affects the out-of-sample performance. To the best of our knowledge, this is the first study to provide a comprehensive analysis of data-driven team selection in FPL, leveraging a hybrid predictive metric within a robust decision-making framework.
  • All datasets, preprocessing scripts, and experiment notebooks are publicly released in a GitHub repository [^3] https://github.com/DanialRamezani/Data-Driven-FPL to ensure full reproducibility and to encourage further research and practical adoption.
@@ 2 Literature review
Several strands of prior work relate to our study [ati2024using]. First, decision-analytic and optimization approaches frame fantasy team management as an explicit choice problem. [matthews2012competing] model the FPL task as a belief-state Markov decision process with a tractable Bayesian Q-learning algorithm, casting weekly choices as a multi-dimensional knapsack and achieving around top-percentile performance against millions of human players. Extending optimization ideas to the NFL, [becker2016analytical] formulate a mixed-integer program to support draft selections and weekly lineup management using historical data and predictions. In FPL, [gupta2019time] propose a linear selection model for squad construction; while intentionally simple and omitting some operational constraints (e.g., optimal formation), they pair it with a hybrid RNN–ARIMA forecaster to supply objective inputs. More recently, metaheuristics and integrated pipelines have appeared: [aribo2024machine] apply a genetic algorithm for team selection and budget allocation, and [venter2024optimisation] combine optimization with forecasting in a practical system that attained a top-$4\%$ rank in the 2021/22 season. Taken together, these studies underscore the promise of explicit decision models, though they differ in how fully they encode FPL rules and whether captaincy is optimized jointly with the starting XI.
@@
A second line of research develops predictive models for player performance, often treating the decision task separately. For NFL quarterbacks, [lutz2015fantasy] evaluated neural networks and support vector machines to improve score prediction, and [steenkiste2015finding] compared linear regression, random forests, and multivariate adaptive regression splines for next-week outcomes. Hybrid time-series ideas enter the literature with [gupta2019time], who fuse a recurrent neural network with ARIMA to estimate the cost vector used downstream. Within FPL, [bangdiwala2022using] assessed several machine-learning models and reported linear regression as the best performer on their dataset, while [pokharel2022fantasy] used an XGBoost regressor to study selection and transfer strategies through a return-on-investment lens and examine the effects of midweek cup fixtures. Deep sequence models have also been explored: [lombu2024predicting] compared convolutional neural networks with long short-term memory networks using form from the previous five matches, finding LSTM superior. Overall, the forecasting literature is methodologically rich, yet downstream optimization is often decoupled from prediction.
@@
Complementing model-based prediction, researchers have examined crowd signals and ancillary tasks that influence managerial choices and market dynamics. [bhatt2019should] analyzed Twitter data for captain selection and found that the “wisdom of the crowd” outperforms expert analyst recommendations in their setting. In a related operational facet, [khamsan2019handling] addressed highly imbalanced classification when predicting virtual player price changes in FPL, where the majority class corresponds to no change.
@@ 3 Preliminaries and mathematical formulations
For simplicity, gamers who select players will be referred to as team owners, while professional Premier League footballers are called players. As previously mentioned, the goal of this paper is to select players in a way that maximizes the total points in the upcoming weeks. Therefore, special and one-time cards are not taken into account. The selection of the starting eleven can be formulated as the following integer programming model:
@@
In this work, various methods are analyzed to estimate the best choice of $c_{j}$ for out-of-sample performance. Although decision-makers are free to define any budget below 100, in this study, we set $b=83.5$. The rationale is straightforward: the model must also account for four reserve players, and in FPL the cheapest players typically cost around four million pounds each, totaling 16 million. An additional 0.5 million is added as a buffer to prevent budget violations. Thus, the formulation focuses on selecting the best fixed starting squad, while team owners are free to choose four inexpensive reserve players for a total of up to £ 16.5 million.
@@ 4 The proposed framework
Finally, a 0–1 integer program selects the starting XI and captain under FPL rules—budget, valid formations, and at most three players per club—with the captain required to start. A bench is then chosen from the remaining budget to complete a 2–5–5–3 squad under the same club quota. Besides maximizing expected points, we also report surrogate XI objectives (xGI-xGC or ICT) and an optional robust variant that penalizes scores by uncertainty ($c_{j}\pm d_{j}$). We output the recommended XI, captain, and bench, along with formation trends, weekly and cumulative points, method leaderboards, and budget/rolling-window sensitivity analyses.
@@ 6 Dataset and experimental setup
Figure 3: Pearson correlation heatmap of modeling features.
Figure 4: RainCloud plots of top numeric features by position.
(a) Goalkeepers
(b) Defenders
(c) Midfielders
(d) Forwards
Figure 5: SHAP value summaries for key features by position.
@@
    • For interpretability, we also fit an XGBoost regressor on the same proxies and use SHAP to visualize influential features. This step is diagnostic only and is not fed to the optimizer in the reported runs.
@@ 8 Experimental results and discussion
(a) Simple average
(b) Weighted average
Figure 11: Budget sensitivity: averaging methods (static teams).
(a) ARIMA (1,0,0) (rolling)
(b) ARIMA (1,0,1) (rolling)
Figure 12: Rolling selection: ARIMA (1,0,0) and ARIMA (1,0,1).
@@
We evaluate models by out-of-sample cumulative points. Training uses gameweeks 1–26 of the 2023/24 season; evaluation spans gameweeks 27–38. In gameweek 29, many player records are missing; instead of discarding teams, we follow the previously explained bench substitution policy. All plots report cumulative team points (no chips). To avoid clutter, we show the best-performing variant within each method family.
Fig. 6 compares averaging-based estimators, and their robust counterparts, to simulation approaches (bootstrap and parametric Monte Carlo, each with B=1000B{=}1000 paths). It can be observed that (i) Weighted average (recency) consistently dominates simple averaging and exponential smoothing, indicating that recent form provides a useful bias signal for FPL points. (ii) Robustification does not help the averaging family: conservatism against point overestimation sacrifices upside without materially reducing downside weeks, so robust curves largely track or underperform their deterministic counterparts. (iii) Bootstrap vs. Monte Carlo are near indistinguishable across the horizon, suggesting that, at this granularity, distributional assumptions add little beyond the empirical resampling of historical points.
Fig. 7 summarizes nine ARIMA specifications. It can be observed that simpler ARIMA structures generalize better: ARIMA (0,0,1) and ARIMA (1,0,1) exceed 600 cumulative points, ARIMA (1,0,0) trails slightly, and performance degrades for more complex models (e.g., ARIMA (2,1,2)). This is consistent with the short, noisy weekly horizon where over-parameterized models overfit in-sample fluctuations that do not persist.
Fig. 8 contrasts ICT-based objectives with an attack-defense surrogate that maximizes expected goal involvement minus expected goals conceded (EGI-EGC). It can be observed that robust ICT modestly improves over deterministic ICT, unlike the averaging family, but the gain is small. The involvement proxy underperforms, likely because conceding risk is not individualized well at the player level and is already indirectly captured by team quality and minutes.
(a) ARIMA (0,0,1)
(b) ARIMA (1,0,0)
(c) ARIMA (1,0,1)
Figure 13: Budget sensitivity: ARIMA family (static teams).
(a) ICT
(b) Hybrid ICT (1:2)
(c) Monte Carlo
Figure 14: Budget sensitivity: ICT objectives and Monte Carlo (static teams).
(a) Simple average (rolling)
(b) Weighted average (rolling)
(c) ARIMA (0,0,1) (rolling)
Figure 15: Rolling selection: averaging family and ARIMA (0,0,1).

### v3

@@ Abstract
Fantasy football is a billion-dollar industry with millions of participants. Under a fixed budget, managers select squads to maximize future Fantasy Premier League (FPL) points. This study formulates lineup selection as data-driven optimization and develops deterministic and robust mixed-integer linear programs that choose the starting eleven, bench, and captain under budget, formation, and club-quota constraints (maximum three players per club). The objective is parameterized by a hybrid scoring metric that combines realized FPL points with predictions from a linear regression model trained on match-performance features identified using exploratory data analysis techniques. The study benchmarks alternative objectives and cost estimators, including simple and recency-weighted averages, exponential smoothing, autoregressive integrated moving average (ARIMA), and Monte Carlo simulation. Experiments on the 2023/24 Premier League season show that ARIMA with a constrained budget and a rolling window yields the most consistent out-of-sample performance; weighted averages and Monte Carlo are also competitive. Robust variants and hybrid scoring metrics improve some objectives but are not uniformly superior. The framework provides transparent decision support for fantasy roster construction and extends to FPL chips, multi-week rolling-horizon transfer planning, and week-by-week dynamic captaincy.
@@ 1 Introduction
Against this backdrop, fantasy football has attracted attention across research and practitioner communities. Qualitative studies examine impacts on related markets and behavior, including merchandise sales (drayer2010effects), game attendance (nesbit2010impact), and participant well-being (wilkins2021exploring). Quantitative research has explored performance prediction and strategy design using econometric and machine learning techniques. The increasing availability of public data and analytics tooling has further lowered barriers to empirical work and large-scale evaluation in this domain.
Despite this progress, most quantitative FPL studies emphasize forecasting weekly player points, with comparatively less attention to integrating those forecasts into a decision model that captures the full set of FPL constraints and jointly selects the starting XI and captain. Decision-analytic approaches exist (matthews2012competing), but captaincy is typically handled outside the optimization pipeline (e.g., via crowds or heuristics) rather than optimized jointly (bhatt2019should). Moreover, robustness to estimation error is rarely addressed, and there is limited comparative evidence on how alternative objective functions, selection budget, rolling-window specifications, or cost-vector estimation methods affect out-of-sample outcomes within a unified optimization framework.
@@
  • We formulate FPL lineup selection as a data-driven optimization problem and develop deterministic and robust integer-programming models that choose the starting XI, bench players, and captain under budget, formation, and club-quota constraints.
  • We propose a hybrid scoring metric that combines realized FPL points with predictions from a ridge regression model trained on interpretable match-performance features.
  • We build a unified pipeline that benchmarks alternative objectives and multiple forecasting approaches (including recency-weighted averaging, low-order ARIMA models, and simulation-based estimators) and provide an empirical assessment on the 2023/24 Premier League season. This analysis includes a rolling-window specification and examines how constraining the budget for the starting XI affects out-of-sample performance, yielding managerial insights into emergent formations and player selection patterns. To the best of our knowledge, this is among the first systematic studies of data-driven team selection in FPL that leverages a hybrid predictive metric within a robust decision-making framework and provides such comprehensive analysis.
  • We publicly release all datasets, preprocessing scripts, and experiment notebooks in a GitHub repository [^3] https://github.com/DanialRamezani/Data-Driven-FPL to facilitate reproducibility and encourage further research and practical adoption.
@@ 2 Related work
Several strands of prior work relate to our study (ati2024using). First, decision-analytic and optimization approaches frame fantasy team management as an explicit choice problem. matthews2012competing model the FPL task as a belief-state Markov decision process with a tractable Bayesian Q-learning algorithm, casting weekly choices as a multi-dimensional knapsack and achieving around top-percentile performance against millions of human players. Extending optimization ideas to the NFL, becker2016analytical formulate a mixed-integer program to support draft selections and weekly lineup management using historical data and predictions. In FPL, gupta2019time propose a linear selection model for squad construction; while intentionally simple and omitting some operational constraints (e.g., optimal formation), they pair it with a hybrid RNN–ARIMA forecaster to supply objective inputs. More recently, metaheuristics and integrated pipelines have appeared: aribo2024machine apply a genetic algorithm for team selection and budget allocation, and venter2024optimisation combine optimization with forecasting in a practical system that attained a top-$4\%$ rank in the 2021/22 season. Taken together, these studies underscore the promise of explicit decision models, though they differ in how fully they encode FPL rules and whether captaincy is optimized jointly with the starting XI.
@@
A second line of research develops predictive models for player performance, often treating the decision task separately. For NFL quarterbacks, lutz2015fantasy evaluated neural networks and support vector machines to improve score prediction, and steenkiste2015finding compared linear regression, random forests, and multivariate adaptive regression splines for next-week outcomes. Hybrid time-series ideas enter the literature with gupta2019time, who fuse a recurrent neural network with ARIMA to estimate the cost vector used downstream. Within FPL, bangdiwala2022using assessed several machine-learning models and reported linear regression as the best performer on their dataset, while pokharel2022fantasy used an XGBoost regressor to study selection and transfer strategies through a return-on-investment lens and examine the effects of midweek cup fixtures. Deep sequence models have also been explored: lombu2024predicting compared convolutional neural networks with long short-term memory networks using form from the previous five matches, finding LSTM superior. Overall, the forecasting literature is methodologically rich, yet downstream optimization is often decoupled from prediction.
@@
Complementing model-based prediction, researchers have examined crowd signals and ancillary tasks that influence managerial choices and market dynamics. bhatt2019should analyzed Twitter data for captain selection and found that the “wisdom of the crowd” outperforms expert analyst recommendations in their setting. In a related operational facet, khamsan2019handling addressed highly imbalanced classification when predicting virtual player price changes in FPL, where the majority class corresponds to no change.
@@ 3 Preliminaries and mathematical formulations
For simplicity, gamers who select football players will be referred to as team owners, while professional Premier League footballers are called players. As previously mentioned, the goal of this paper is to select players in a way that maximizes the total points in the upcoming weeks. Therefore, special and one-time cards are not taken into account. The selection of the starting eleven can be formulated as the following integer programming model:
@@
In this work, various methods are analyzed to estimate the best choice of $c_{j}$ for out-of-sample performance. Although decision-makers are free to define any budget below 100, in this study, we set $b=83.5$ as a default value. The rationale is straightforward: the model must also account for four reserve players, and in FPL the cheapest players typically cost around four million pounds each, totaling 16 million. An additional 0.5 million is added as a buffer to prevent budget violations. Thus, the formulation focuses on selecting the best fixed starting squad, while team owners are free to choose four inexpensive reserve players for a total of up to £ 16.5 million.
@@ 4 The proposed framework
Finally, a 0–1 integer program selects the starting XI and captain under FPL rules-budget, valid formations, and at most three players per club-with the captain required to start. A bench is then chosen from the remaining budget to complete a 2–5–5–3 squad under the same club quota. Besides maximizing expected points, we also report surrogate XI objectives (xGI-xGC or ICT) and an optional robust variant that penalizes scores by uncertainty ($c_{j}\pm d_{j}$). We output the recommended XI, captain, and bench, along with formation trends, weekly and cumulative points, method leaderboards, and budget/rolling-window sensitivity analyses.
@@ 6 Dataset and experimental setup
    • For interpretability, we also fit an XGBoost regressor on the same proxies and use SHAP library to visualize influential features. This step is diagnostic only and is not fed to the optimizer in the reported runs.
@@ 7 Exploratory data analysis
Figure 3: Pearson correlation heatmap of modeling features.
Figure 4: RainCloud plots of top numeric features by position.
(a) Goalkeepers
(b) Defenders
(c) Midfielders
(d) Forwards
Figure 5: SHAP value summaries for key features by position.
Figure 6: Hierarchical clustering of strategy behaviors. The heatmap reports the absolute Spearman correlation between weekly team scores for all strategies, and the dendrogram groups runs with similar week-to-week profiles.
@@ 8 Experimental results and discussion
Figure 6 summarizes how all strategies move week-to-week. Each cell is the absolute Spearman correlation between two methods’ weekly team scores (GW27-38), and the dendrogram clusters methods with similar temporal profiles. Clear block-diagonal structure appears: rolling ARIMA variants form a tight cluster, simple/weighted/ExpSm averages cluster together, simulation/Monte-Carlo runs form another block, and ICT/linear-regression-based strategies sit farther away with weaker affinities to the averaging/ARIMA families. Budgets of the same method line up next to one another (e.g., £80-£65), indicating that the budget mainly shifts level but not the week-to-week pattern. Hybrid methods typically lie between their parent families (e.g., Hybrid Weighted near Weighted Avg.), while robust versions are slightly less correlated, reflecting their penalty on uncertainty. Overall, the map shows substantial redundancy—many methods behave very similarly—so in the remainder we focus on a representative subset and benchmark uplifts relative to a single strong baseline.
@@
Having analyzed the similarity of strategy behaviors (Figure 3), we now focus on the models that achieve the strongest out-of-sample performance. We evaluate strategies by cumulative team points, training on gameweeks 1-26 of the 2023/24 season and evaluating on gameweeks 27-38. In gameweek 29 many player records are missing; instead of discarding teams, we apply the bench-substitution policy described earlier. All subsequent plots report cumulative team points (with no chips played). To avoid clutter and highlight practical takeaways, we show only the best-performing variant within each method family in the remaining figures and tables.
Figure 7 reports cumulative out-of-sample points for the averaging and simulation families. Panel (a) shows that the Weighted Avg. (recency-weighted mean) clearly dominates Simple Avg. and Exponential Smoothing over GW 27–38, suggesting that putting more weight on recent form is beneficial for FPL scoring. The robust versions of these methods track their non-robust counterparts from below: penalizing uncertainty mainly trims upside weeks without delivering visible protection in bad weeks, so robust curves never catch up. Panel (b) compares Simulation (Non Parametric) (bootstrap) and Monte Carlo Simulation to the simple average baseline; both simulation curves almost coincide with each other and with Simple Avg., indicating that, at this horizon and sample size, adding a stochastic layer (either empirical or parametric) offers little advantage over a deterministic average of past points.
Figure 8 compares nine ARIMA specifications. The simplest models without differencing, ARIMA(0,0,1) and ARIMA(1,0,1), clearly dominate in this static setting, finishing the horizon at roughly 680 and 660 cumulative points, respectively. ARIMA(1,0,0) forms a second tier, ending about 50 points behind, while all higher-order or differenced variants (e.g., ARIMA(1,1,1), ARIMA(2,1,1), ARIMA(2,1,2)) lag steadily throughout GW 27–38. This pattern is consistent with the short, noisy weekly horizon: lightly parameterized ARIMA models capture the main signal, whereas more complex configurations tend to overfit in-sample fluctuations that do not persist out of sample. However, when ARIMA processes are embedded in our rolling-selection framework with a constrained starting-XI budget, ARIMA(1,0,0) proves more stable and ultimately delivers the best overall performance; for this reason, ARIMA(1,0,0) with a £70m XI budget is adopted as the benchmark ARIMA model in our later cross-family and external comparisons.
Figure 9 compares alternative objectives: ICT, robust ICT, and an attack-defense proxy based on maximizing expected goal involvement minus expected goals conceded (EGI-EGC). It is evident that robust ICT marginally outperforms deterministic ICT, continuing the trend seen in the averaging family, though the improvement is modest. On the other hand, the involvement proxy underperforms significantly, likely due to the inability to capture indirect or additional important features such as those related to team quality, minutes played, and so on.
We next study the effect of the hybrid scoring metric in Eq. (28). Figure 10 applies it to the averaging and simulation families, while Figure 11 applies it to ARIMA, ICT, and linear regression. Across all families, the 1:2 setting (placing two-thirds of the weight on the model-based score and one-third on realised FPL points) is the most reliable: it typically sits above or on top of the original method’s curve, delivering a modest but consistent uplift. In contrast, the 2:1 setting often underperforms the base method, suggesting that overweighting raw historical points re-introduces the very noise the predictive model is designed to filter out.
For the stronger baselines—ARIMA (0,0,1), ICT, and linear regression—the same pattern holds: the hybrid 1:2 variants track or exceed the non-hybrid trajectories over most of the horizon, whereas the 2:1 versions are uniformly worse. This indicates that hybridization is most effective when it uses realized points as a light calibration signal on top of a stable predictive score, rather than as the dominant driver of team selection.
Figure 12: Budget sensitivity (static teams): Simple average.
The base budget is our default b = £83.5 for the starting XI (with the total FPL budget still at £100). We then vary this cap between £55 and £80 to study sensitivity. Figure 12 shows how different budget caps affect a single static team built with the simple‐average predictor. The bump chart tracks the rank (1 = best) of each budget’s cumulative score from GW 27 to 38, with final totals reported on the right. After some early crossings, the ranking stabilises: the £80 budget quickly takes over and remains first to the end, followed by £75, while the nominal “Base” budget finishes only fourth behind £70. The two lowest caps (£60 and especially £55) stay near the bottom throughout. Overall, the figure suggests clear but non-linear budget sensitivity: moving from very low to moderate budgets yields large gains, but simply increasing the cap beyond about £75 does not guarantee better static performance.
Figure 13: Budget sensitivity (static teams): Weighted average.
Figure 13 illustrates the budget sensitivity for static teams using the weighted average method. The plot shows how teams with different budgets (ranging from £80 to £60) perform across the gameweeks. It is evident that teams with a higher budget consistently rank better, with the £80 budget achieving the highest rank and maintaining strong performance throughout the season. Conversely, teams with lower budgets, particularly £60, underperform, indicating that a more substantial budget enables better team selection and performance. This reinforces the importance of budget flexibility in achieving competitive rankings, as the weighted average method benefits from selecting players with higher recent performance, which requires greater financial resources.
Figure 14: Rolling selection: ARIMA (1,0,0).
Figure 14 illustrates the budget sensitivity and performance of static and rolling selection strategies using ARIMA (1,0,0). The plot compares static teams with varying budgets (£80, £75, £70, £65, £60) and rolling teams for each corresponding budget. While static teams with higher budgets generally perform better, the rolling selection strategy (denoted by the "Roll" lines) exhibits competitive performance, especially for the £75 and £70 budgets, which outperform the static £80 variant during some gameweeks. This highlights that rolling selection can adapt dynamically to ongoing changes and might offer an edge over static selection in certain budget ranges, even if it has higher variance in rankings.
Figure 15: Rolling selection: ARIMA (1,0,1).
Figure 15 shows the budget–rank trajectories for static and rolling teams under ARIMA (1,0,1). Rolling selection is again highly competitive: the £80 and £70 rolling variants dominate the ranking from roughly GW 31 onward and finish the season as the top two configurations. The static base and £80–£75 teams form a second tier, while low-budget static teams (£65–£60) remain near the bottom throughout. The £60 rolling team still outperforms its static £60 counterpart, indicating that dynamic re-optimization helps even when resources are tight, although the benefit is most pronounced at higher budgets.
Figure 16: Budget sensitivity (static teams): ARIMA (0,0,1).
Figure 16 summarizes budget sensitivity for static ARIMA (0,0,1) teams. The coloured strip at the top shows which budget produced the highest weekly score in each gameweek (GW 27–38), while the lollipops on the bottom panel report the GW 38 scores for each budget together with the number of weekly “wins” (W) in the legend. The £75 variant records the most weekly wins (4), followed by the base-budget squad (3) and the £80 and £65 variants (2 each), indicating that mid-range budgets frequently top the table in individual weeks. However, the base team still attains the highest final-week score (41 points), with £75 close behind and the low-budget £65 side clearly lagging. Overall, ARIMA (0,0,1) exhibits only modest sensitivity to budget: extra funds offer some benefit at the extremes, but the performance gap between the base and mid-range budgets is small relative to week-to-week volatility.
Figure 17: Budget sensitivity (static teams): ARIMA (1,0,0).
Figure 17 reports budget sensitivity for static ARIMA (1,0,0) teams. The top strip shows which budget wins each gameweek; most tiles are orange or green, confirming that the £80 and £75 variants dominate weekly leadership (four wins each), while the base and cheaper budgets only top the table once. The lollipops in the lower panel summarize GW 38 performance: the £80 squad achieves the highest final score (49 points), followed by £75 (43) and the base budget (39), with £70, £60, and £65 trailing more clearly. Overall, ARIMA (1,0,0) benefits more from extra budget than ARIMA (0,0,1): mid-to-high budgets (£75–£80) deliver both more weekly wins and a noticeable improvement in final-week performance over the base case.
Figure 18: Budget sensitivity (static teams): ARIMA (1,0,1).
Figure 18 shows budget sensitivity for static ARIMA (1,0,1) teams. The top winner strip indicates which budget leads in each gameweek: the £75 configuration is the most frequent weekly leader (five wins), while the base, £70, and £60 budgets each top the table in only two weeks, and £80 wins once; £65 never leads. The lollipop plot summarizes performance in the final gameweek: the £80 and base squads achieve the highest totals (55 and 54 points), followed by £75 (48 points), whereas the cheaper budgets (£70, £65, £60) cluster around the mid-20s. Overall, ARIMA (1,0,1) still rewards higher budgets, but the £75 squad offers a good compromise, capturing most weekly wins while using less budget than the full £80 baseline.
Figure 19: Budget sensitivity (static teams): ICT objective.
Figure 20: Budget sensitivity (static teams): Hybrid ICT (1:2).
Figure 19 reports budget sensitivity for the ICT-based objective with static squads. The winner strip along the top shows which budget dominates each gameweek: the base configuration wins most often (5 weeks), followed by the £80 and £70 squads (3 and 2 wins, respectively), while the £75 team wins in two late gameweeks. The lollipop panel summarizes season-end performance. Here, the £75 team achieves the highest final weekly total ($\approx$ 59 points), with the base and £80 squads essentially tied just below 50 points, and the £70 budget clearly lagging (46 points). Thus, for ICT the mid-range £75 budget offers the best overall outcome, while spending up to £80 yields little extra value and cutting to £70 is noticeably harmful.
Figure 20 shows budget sensitivity for the Hybrid ICT (1:2) objective with static squads. The winner strip indicates that leadership is more evenly shared than for pure ICT: the £80 and £70 squads each top several gameweeks (three apiece), while the base, £75, and £65 teams each lead twice. The lollipop panel summarizes the final weekly totals. Here the base-budget hybrid team again finishes first ($\approx$43 points), followed by £80 and £70 ($\approx$38 and 36 points), with £65 and £75 clearly trailing in the mid-20s. Thus, under Hybrid ICT, modest budget increases above the baseline offer only limited gains, whereas pushing the budget down to £65–£75 substantially weakens the squad, suggesting that the hybrid objective extracts most of its value from model structure rather than additional spending.
Figure 21: Budget sensitivity (static teams): Monte Carlo.
Figure 22: Rolling selection: Simple average.
Figure 21 shows how the Monte Carlo–based static teams react to budget changes. The winner strip indicates that the full-budget baseline and the £80 variant share most of the weekly wins, with the £75 team occasionally coming out on top and the cheaper £70–£60 squads almost never leading in any gameweek. The lollipop plot confirms this pattern: the baseline and £80 teams finish with very similar final weekly totals, £75 is moderately behind, and performance drops sharply once the budget falls to £70 or below. Overall, the Monte Carlo objective benefits from reasonable spending but is relatively flat between the base and £80 budgets, while aggressive under-spending severely limits its upside.
Figure 22 summarizes how the rolling Simple Avg. strategy reacts to budget changes. The winner strip shows that the £80 variant wins the largest number of individual gameweeks (five), followed by the baseline budget (four), while the £70 side wins only three weeks and the £65 team never tops the weekly ranking. The lollipop plot at the bottom reinforces this ordering: the £80 squad achieves the highest final weekly total, the baseline team is a close second, and performance drops sharply once the budget is cut to £70–£65. Overall, the rolling simple-average strategy benefits from modest extra spending but is quite fragile under tighter budget constraints.
Figure 23: Rolling selection: Weighted average.
Figure 24: Rolling selection: ARIMA (0,0,1).
Figure 23 shows budget sensitivity for the rolling Weighted Avg. strategy. The winner strip at the top indicates that no single budget dominates week by week: the £80 and £60 variants each top four and three gameweeks respectively, while the baseline budget also wins three weeks and the £70 side only two. However, the lollipop chart reveals that the baseline squad ultimately converts these weekly results into the strongest season total (42 points in the final week), with the £80 team slightly behind and the £70 and £60 teams trailing more clearly. Thus, under rolling re-optimization, the recency-weighted strategy is fairly competitive across budgets, but deep cuts to £60 still incur a noticeable loss in final performance.
Figure 24 (rolling selection with ARIMA(0,0,1)) shows that the rolling ARIMA variant again favors less-constrained budgets: the baseline squad attains the highest final weekly total (annotated on the plot $\approx$41), followed by the £70 team ($\approx$34). The £80 and £60 budgets finish noticeably lower ($\approx$26 and $\approx$25 respectively). The thin winner–strip above the lollipop shows that leadership still switches between budgets week-to-week rather than being monopolized by one budget, but those weekly wins translate unevenly into season-end totals — the baseline configuration converts transient advantages into the largest aggregate return. In short, ARIMA(0,0,1) with rolling re-optimization produces the strongest final outcome for the baseline setup, while tightening the budget (£60) materially reduces end-of-season payoff; intermediate budgets (£70/£80) perform between these extremes.